Solve et + 2e-3 – 3= 0 algebraically.
Answer #1

+22n 3=0 - ex+ 2 -3=0 ex Cex²_3e²+2.0 (ex) 2 2e *_ ex+2=0 - excex-2)-1 (0-2)=0 (ex-) (ex-2)=0 27 ex = 2 - lage - heg In (1) I